OpenStudy (anonymous):

3. In a city council election, 2445 votes were cast for one of the candidates. If the candidate received at least one-third of the votes, how many voters, voted in the election? a. write an inequality to represent this situation, using x to represent the number of voters that voted in the election. b. Solve the inequality by isolating the variable. Show all work leading to your answer. c. Write 1-2 sentences to explain what the solution to the inequality represents for this situation.

OpenStudy (texaschic101):

(1/3)x > = 2445 (thats greater then or equal) x >= 2445/(1/3) x >= 2445 * 3/1 x >= 7335 The number of voters (x) is greater then or equal to 7335.